# StereoDiffusion: Training-Free Stereo Image Generation Using Latent Diffusion Models

### v1

@@ Abstract
The demand for stereo images increases as manufacturers launch more XR devices. To meet this demand, we introduce StereoDiffusion, a method that, unlike traditional inpainting pipelines, is trainning free, remarkably straightforward to use, and it seamlessly integrates into the original Stable Diffusion model. Our method modifies the latent variable to provide an end-to-end, lightweight capability for fast generation of stereo image pairs, without the need for fine-tuning model weights or any post-processing of images. Using the original input to generate a left image and estimate a disparity map for it, we generate the latent vector for the right image through Stereo Pixel Shift operations, complemented by Symmetric Pixel Shift Masking Denoise and Self-Attention Layers Modification methods to align the right-side image with the left-side image. Moreover, our proposed method maintains a high standard of image quality throughout the stereo generation process, achieving state-of-the-art scores in various quantitative evaluations. The code is available here.
Keywords XR, Deep Image/Video Synthesis, Image Editing, Artificial Intelligence, Inpainting, Stable Diffusion
@@ 1 Introduction
Figure 1: Our approach takes one of three types of user input and generates a stereo image. The accepted user inputs are (a) a photo, (b) a text prompt, or (c) a user’s image as a depth map and a prompt. We use a latent diffusion model pretrained on images for inputs a and b and on depth maps for input c.
@@
Large-scale language-image (LLI) models have become prominent in recent years, acclaimed for their advanced generative semantic and compositional abilities [1, 2, 3, 4, 5]. Their distinctiveness lies in their training on extensive language-image datasets, enabling them to interpret and generate content from diverse linguistic and visual contexts. Utilizing innovative image generative techniques such as auto-regressive and diffusion models [6], these LLI models have significantly advanced the synergy between linguistic understanding and image generation. This has led to a new era in creative and semantically rich image synthesis, marking a notable advancement in artificial intelligence and computer vision.
@@
A significant recent development in the VR/AR field is Apple’s introduction of Vision Pro, which has the potential to drive rapid advancements in this field. Despite the growing production of 3D content by various manufacturers and related research [7, 8, 9, 10, 11] in recent years, the availability of stereo multimedia content, which offers a depth-enhanced visual experience, remains relatively scarce. As the VR/AR era looms, the limitations of existing image generation models that are confined to producing 2D images become increasingly apparent. However, there is currently no relevant research that attempts to use image generation models to directly generate stereo image pairs. In response to this challenge, we introduce a novel methodology. Through modification of the Stable Diffusion model’s latent variable, we have devised an efficient end-to-end approach, eliminating the need for additional models like inpainting [12, 13] for post-processing to generate stereo images. Examples in Fig. 1. We address the constraints of traditional image generation models that employ an inpainting pipeline. Our approach is to generate stereo image pairs by adjusting the latent variable of the Stable Diffusion model, see Fig. 2. We utilize Symmetric Pixel Shift Masking Denoise and Self-Attention layers modification to align the generated right-side image with the left-side image. This method allows for a lightweight, fine-tune free solution that can be seamlessly integrated into the original Stable Diffusion model without the need for model fine-tuning. To the best of our knowledge, our approach represents the first instance of generating stereo images by modifying the latent variable of Stable Diffusion. Compared with other methods, our approach enables the training-free end-to-end rapid generation of high-quality stereo images using only the original Stable Diffusion model.
@@
Figure 2: Pipeline of Stereo Diffusion. The pipeline illustrates the process of starting with random noise and denoising it to generate stereo image pairs. The operation of Stereo Pixel Shift is represented by Eq. 3. The Disparity Map for generating stereo image pairs can be obtained from depth models such as DPT [14] or MiDas [15]. The pipeline only shows the Unidirectional Self-Attention operation, designed to align the right-side image with the left-side image, a method that satisfies general needs. Bidirectional Self-Attention, being a mutual operation, would be represented by bidirectional arrows in the image. The orange box in the image depicts the concept of Symmetric Pixel Shift Masking Denoise, with details explained in Section 3.2. The cross attention part of the sampling process is omitted for brevity.
@@ 2 Related work
Diffusion models, notably the Denoising Diffusion Implicit Models (DDIM) [16], have made significant strides in image generation. The DDIM sampling algorithm revealed that using the same initial noise results in consistent high-level features across different generative paths, indicating initial noise as a potent latent image encoding [16]. This discovery aids in modifying images by adjusting the Stable Diffusion latent variable. A key challenge in stereo image generation is maintaining content consistency between paired images. Researchers are focusing on image editing techniques using Stable Diffusion [17, 18, 19, 20], such as the "prompt-to-prompt" method [6], which involves altering the model’s cross-attention during sampling for text-prompt-based image editing. ControlNet [21] is also a notable work in the field, but instead of utilizing the latent space, the authors trained ControlNet on a large dataset to better control the generation of desired images by Stable Diffusion. Additionally, ControlNet primarily focuses on pose control and lacks the capability for pixel-level modifications of images. Although effective, these methods are less suited for tasks needing precise pixel-level manipulation, like stereo image generation, due to their reliance on text prompts for image modification.
@@
Ensuring the consistency of images within the same batch in Stable Diffusion has long been a challenge in video generation [22, 23]. VideoComposer addressed this by incorporating an STC-encoder into the Latent Diffusion Model’s U-Net model, ensuring consistency in the generated image content [22]. Similarly, VideoLDM achieved impressive video generation results by introducing 3D convolution layers and temporal attention layers into the spatial and temporal layers of U-Net [23]. However, these methods require fine-tuning of the original models and substantial amounts of data. Generally, this is not an issue for video generation, but for aiming to achieve stereo image generation, the available stereo image data is quite limited, mostly comprising road traffic images initially intended for autonomous driving depth prediction services. There have been attempts to explored zero-shot video generation in the video generation field [24, 25], such as Tune-A-Video [24]. This work utilized a technique called ST-Attn to maintain the continuity of videos. We employ a comparable approach to ensure consistency between the left and right images.
@@
Traditional image-based reconstruction and rendering methods require complex capture setups, involving numerous images with significant baselines[26, 27, 28, 29]. Currently, there are limited research endeavors directly focused on generating stereo images. Many studies have concentrated on generating 3D photos, a technique allowing subtle changes in the camera angle for observing photos from different perspectives [13, 12, 30, 31, 32]. Among 3D image generation techniques, 3D Photography Inpainting is a notable approach [12, 26]. This method employs inpainting to generate 3D images. After passing the input image through a depth estimation model, they map the image onto a mesh and apply changes in perspective based on the depth map of the original image. Inpainting is then utilized to fill the gaps left by transformed pixels in the original image. This approach significantly differs from our modification of the Stable Diffusion latent space. Although this method could be adapted as post-process after image generated through Stable Diffusion, it requires additional steps and consumes more time.
3D scene generation by pretrained Stable Diffusion
Recently, numerous studies have employed model distillation techniques using the 2D image priors of pre-trained Stable Diffusion models for text-based 3D model reconstruction. A notable work in this field is Dreamfusion [33], where researchers utilized a method known as ’Score Distillation Sampling’ (SDS). This method involves initializing a NeRF-like model with random weights and repeatedly rendering views of this NeRF from random camera positions and angles. These renderings are then used as inputs for an Imagen-surrounding score distillation loss function. Subsequently, other researchers improved upon SDS, proposing Variational Score Distillation (VSD) [34], which significantly enhances the quality of generated 3D scenes. In theory, these methods can be used to create 3D scenes and then produce stereo image pairs using rendering-based techniques. However, currently, these methods require several hours to generate complete 3D scenes. Some users might only need simple stereo image pairs, not the entire 3D scene, necessitating a lightweight, rapid method for generating stereo image pairs. Our method offers a fast, end-to-end solution for creating stereo image pairs.
3 Methods
@@
Diverging from conventional inpainting methods, our approach is distinctively simple and training-free. It seamlessly integrates into the original Stable Diffusion framework for end-to-end generation of stereo image pairs, eliminating the need for post-processing. Our method leverages a disparity map in the early denoising stage to apply a Stereo Pixel Shift (Section 3.1) to the latent vector of the left image. This process generates the latent vector for the right image through disparity. To address the inconsistency issues between the left and right images during the denoising process, we employ a Symmetric Pixel Shift Masking Denoise (Section 3.2) technique and a Self-Attention module (Section 3.3) to align the right image with the left one. Since our method exclusively manipulates the latent variable, it can be applied across various image generation tasks in different Stable Diffusion models. This versatility stems from the technique’s focus on latent space operations, making it adaptable to a wide range of scenarios within the Stable Diffusion framework (Section 3.4). Our method only requires a disparity map which can be obtained by various depth estimation models like DPT [14], MiDas [15] etc. and does not require camera calibration.
@@ 3.1 Stereo Pixels Shift
For the task of generating stereo images, fine-tuning models on large stereo datasets like KITTI seems intuitive. However, after fine-tuning the model using various methods such as ControlNet [35] and Lora [36], the results of the generated images remains unsatisfactory. A major flaw of this approach is that even if we could generate high-quality stereo image pairs, the types of images generated will be limited to driving scenes similar to KITTI, losing the most important feature of Stable Diffusion: its diversity. Inspired by the Denoising Diffusion Implicit Models (DDIM) sampling technique for Stable Diffusion [16], we present a new method, Stereo Pixels Shift, without the aforementioned drawbacks. Utilizing DDIM for sampling from generalized generative processes, a latent vector sample $\bm{x}_{t-1}$ is generated from a sample $\bm{x}_{t}$ via a noise predictor $\epsilon_{\theta}$:
$\bm{x}_{t-1}=\>$ $\sqrt{\alpha_{t-1}}\,\underbrace{\left(\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}^{(t)}\!\!\left(\bm{x}_{t}\right)}{\sqrt{\alpha_{t}}}\right)}_{\text{predicted }\bm{x}_{0}}$ (1)
@@
where $\epsilon_{t}$ is noise following a standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\bm{x}_{t}$, and $\alpha_{t}$ controls the noise scale at step $t$ with $\alpha_{0}:=1$. If we set $\sigma_{t}=0$ for all $t$ and the same model $\epsilon_{\theta}$ is used, the generative results are consistent and identical, making the forward process deterministic, given $\bm{x}_{t-1}$ and $\bm{x}_{0}$. Thus the result of $\bm{x}_{t-1}$ depends solely on $\bm{x}_{t}$. During the denoise process at a certain step $t^{\prime}$, if we modify $\bm{x}_{t^{\prime}}$ to $\bm{x}^{\prime}_{t^{\prime}}$, subsequently, $\bm{x}^{\prime}_{t^{\prime}-1}$ is denoised based on $\bm{x}^{\prime}_{t^{\prime}}$, eventually generating $\bm{x}^{\prime}_{0}$ which is different from the original $\bm{x}_{0}$. This pivotal insight enables the practical application of Stable Diffusion for stereo image generation. To align with this approach, we scale down the disparity map to match the dimensions of the latent space. Subsequently, we manipulate the latent vector on a pixel-by-pixel basis, guided by the disparity map. Given the relatively small size of the latent vector, this process does not entail a substantial computational overhead.
Assuming that the two images have parallel optical axes, disparity maps can be derived from depth maps based on
@@
The Stereo Pixel Shift operation $\mathcal{S}$ can be expressed as
@@
The reason that we can apply Stereo Pixel Shift to latent variable is that, after a certain step, there is a spatial position correspondence between the latent variable and the generated image. According to diffusion process theory [38, 39, 16, 20, 40, 41], sampling can be represented as
@@ 3.2 Symmetric Pixel Shift Masking Denoise
According to Eq. 1, a pixel shift applying to $x_{t^{\prime}}$ to obtain $x^{\prime}_{t^{\prime}}$ results in a slight difference between $x^{\prime}_{t^{\prime}-1}$ and $x_{t^{\prime}-1}$. This difference accumulates during the subsequent denoising process, leading to variations in the final generated image. As a result, when the denoising algorithm is applied, it may interpret the shifted areas differently, potentially causing variations in how the subject matter appears after processing. This challenge is crucial in stereo image generation, as maintaining symmetry and coherence between the two sides is essential for creating a convincing and realistic stereo effect.
To circumvent the issue, inspired by the concept of inpainting, we propose the Symmetric Pixel Shift Masking Denoise method. We create a mask for the area where the stereo pixel shift is applied. At regular intervals, defined by specific steps $t^{\prime}$, the values from the masked region of the left latent space are copied to the corresponding area of the mask in the right latent space. Consequently, the denoising process for the right image can be reformulated from Eq. 1 as
$\bm{x}^{\prime}_{t^{\prime}-1}=\>$ $\sqrt{\alpha_{t^{\prime}-1}}\,{\left(\frac{\bm{X}^{\prime}_{t^{\prime}}-\sqrt{1-\alpha_{t^{\prime}}}\,\epsilon_{\theta}^{(t^{\prime})}\!\left(\bm{x}^{\prime}_{t^{\prime}}\right)}{\sqrt{\alpha_{t^{\prime}}}}\right)}$ (6)
@@
where $\mathcal{S}$ represents the operation of Stereo Pixel Shift in Eq. 3, $D$ is the corresponding disparity map of the image, and $\mathbf{M}$ is a Boolean matrix of the same shape as $\bm{x}$ that signifies the mask, with values set to True for the pixels that have been shifted. The variable $\bm{x}_{t^{\prime}-1}$ denotes the latent vector of the left image at timestep $t^{\prime}-1$, which we represent by
$\bm{x}_{t^{\prime}-1}=\>$ $\sqrt{\alpha_{t^{\prime}-1}}\,{\left(\frac{\bm{x}_{t^{\prime}}-\sqrt{1-\alpha_{t^{\prime}}}\,\epsilon_{\theta}^{(t^{\prime})}\!\left(\bm{x}_{t^{\prime}}\right)}{\sqrt{\alpha_{t^{\prime}}}}\right)}$ (8)
$+\sqrt{1-\alpha_{t^{\prime}-1}}\,\epsilon_{\theta}^{(t^{\prime})}\!\left(\bm{x}_{t^{\prime}}\right).$
This is derived from Eq. 1 by setting $\sigma_{t}=0$.
@@
where $\epsilon_{t^{\prime}}$ denotes random noise, $\mathbf{M}$ is the mask same as the one in Eq. 7. However, the effectiveness varies with different images. Sometimes, it may even lead to a decrease in the quality of the generated images. A detailed effects analysis of the Deblur technique is presented in the ablation studies described in Section 4.3.

### v2

@@ Abstract
The demand for stereo images increases as manufacturers launch more extended reality (XR) devices. To meet this demand, we introduce StereoDiffusion, a method that, unlike traditional inpainting pipelines, is training-free and straightforward to use with seamless integration into the original Stable Diffusion model. Our method modifies the latent variable to provide an end-to-end, lightweight method for fast generation of stereo image pairs, without the need for fine-tuning model weights or any post-processing of images. Using the original input to generate a left image and estimate a disparity map for it, we generate the latent vector for the right image through Stereo Pixel Shift operations, complemented by Symmetric Pixel Shift Masking Denoise and Self-Attention Layer Modifications to align the right-side image with the left-side image. Moreover, our proposed method maintains a high standard of image quality throughout the stereo generation process, achieving state-of-the-art scores in various quantitative evaluations.
Figure 1: Our method takes one of three types of user input and generates a stereo image. Accepted user inputs: (a) a photo, (b) a text prompt, or (c) a user’s image as a depth map and a prompt. We use a latent diffusion model pretrained on either images (a, b) or depth maps (c).
@@ 1 Introduction
Large-scale language-image (LLI) models have become prominent in recent years, acclaimed for their advanced generative semantic and compositional abilities [29, 25, 45, 21, 6]. Their distinctiveness lies in their training on extensive language-image datasets, enabling them to interpret and generate content from diverse linguistic and visual contexts. Utilizing innovative image generative techniques such as auto-regressive and diffusion models [10] have significantly advanced the synergy between linguistic understanding and image generation. This has led to a new era of creative and semantically rich image synthesis, marking notable advancements in artificial intelligence and computer vision.
@@
A significant recent development in the VR/AR field is Apple’s introduction of Vision Pro, which has the potential to drive rapid advancements in this field. Despite the growing production of 3D content by various manufacturers, and related research [15, 41, 14, 43, 28] in recent years, the availability of stereo multimedia content, which offers a depth-enhanced visual experience, remains relatively scarce. As the VR/AR era looms, the limitations of existing image generation models that are confined to producing 2D images become increasingly apparent. However, there is currently no relevant research that attempts to use image generation models to directly generate stereo image pairs. In response to this challenge, we introduce a novel methodology. Through modification of the Stable Diffusion model’s latent variable, we have devised an efficient end-to-end approach, eliminating the need for additional models like inpainting [31, 12] for post-processing to generate stereo images. Figure 1 presents some examples. Code is available at https://github.com/lez-s/StereoDiffusion.
We address the constraints of traditional image generation models that employ an inpainting pipeline. Our approach is to generate stereo image pairs by adjusting the latent variable of the Stable Diffusion model, see Figure 2. We use Symmetric Pixel Shift Masking Denoise and Self-Attention layer modifications to align the generated right-side image with the left-side image. This method allows for a lightweight solution that can be seamlessly integrated into the original Stable Diffusion model without the need for fine-tuning. To the best of our knowledge, our approach represents the first instance of generating stereo images by modifying the latent variable of Stable Diffusion. Compared with other methods, our approach enables the training-free end-to-end rapid generation of high-quality stereo images using only the original Stable Diffusion model.
@@
Figure 2: The pipeline of our Stereo Diffusion. The process starts with random noise and denoising of it to generate a stereo image pair. The operation of Stereo Pixel Shift is represented by Eq. 3. The Disparity Map for generating stereo image pairs can be obtained from depth models such as DPT [26] or MiDas [27]. The pipeline only shows the Unidirectional Self-Attention operation, designed to align the right-side image with the left-side image, a method that satisfies general needs. Bidirectional Self-Attention, being a mutual operation, would be represented by bidirectional arrows in the image. The orange box in the image depicts the concept of Symmetric Pixel Shift Masking Denoise, with details explained in Sec. 3.2. The cross attention part of the sampling process is omitted for brevity.
@@ 2 Related work
Diffusion models, notably the Denoising Diffusion Implicit Models (DDIM) [32], have made significant strides in image generation. The DDIM sampling algorithm revealed that using the same initial noise results in consistent high-level features across different generative paths, highlighting initial noise as a potent latent image encoding [32]. This discovery aids in modifying images by adjusting the Stable Diffusion latent variable. A key challenge in stereo image generation is maintaining content consistency between paired images. Researchers are focusing on image editing techniques using Stable Diffusion [3, 35, 22, 44], such as the "prompt-to-prompt" method [10], which involves altering the model’s cross-attention during sampling for text-prompt-based image editing. ControlNet [46] is also a notable work in the field, but instead of utilizing the latent space, the authors trained ControlNet on a large dataset to better control the generation of desired images by Stable Diffusion. Additionally, ControlNet primarily focuses on pose control and lacks the capability for pixel-level modifications of images. Although effective, these methods are less suited for tasks needing precise pixel-level manipulation, like stereo image generation, due to their reliance on text prompts for image modification.
@@
Ensuring the consistency of images within the same batch in Stable Diffusion has long been a challenge in video generation [37, 2]. VideoComposer addresses this by incorporating an STC-encoder into the Latent Diffusion Model’s U-Net architecture, ensuring consistency in the generated image content [37]. Similarly, VideoLDM achieves impressive video generation results by introducing 3D convolution layers and temporal attention layers into the spatial and temporal layers of U-Net [2]. However, these methods require fine-tuning of the original models and substantial amounts of data. Generally, this is not an issue for video generation, but for achieving stereo image generation, the available stereo image data is quite limited, mostly comprising road traffic images initially intended for autonomous driving depth prediction services. In the video generation field, attempts such as Tune-A-Video [42] have explored zero-shot video generation [42, 36]. This work utilizes a technique called ST-Attn to maintain the continuity of videos. We employ a comparable approach to ensure consistency between the left and right images.
@@
Traditional image-based reconstruction and rendering methods require complex capture setups, involving numerous images with significant baselines [8, 40, 9, 13]. Currently, limited research endeavors directly focus on generating stereo images. Many studies have concentrated on generating 3D photos, a technique allowing subtle changes in the camera angle for observing photos from different perspectives [12, 31, 47, 34, 18]. Among 3D image generation techniques, 3D Photography Inpainting is a notable approach [31, 8]. This method employs inpainting to generate 3D images. After passing the input image through a depth estimation model, they map the image onto a mesh and apply changes in perspective based on the depth map of the original image. Inpainting is then utilized to fill the gaps left by transformed pixels in the original image. This approach significantly differs from our modification of the Stable Diffusion latent space. Although this method could be adopted as post-processing of an image generated through Stable Diffusion, it requires additional steps and consumes more time.
3D scene generation by pretrained Stable Diffusion.
Recent studies have used model distillation with pretrained Stable Diffusion models for text-based 3D reconstruction. DreamFusion [23] employes ‘Score Distillation Sampling’ (SDS) to initialize and render a NeRF model, improving with Imagen-surrounding score distillation loss. Variational Score Distillation (VSD) [38] further enhances 3D scene quality. These methods can create stereo images via rendering but are time-intensive. Without full 3D scenes, our method provides a faster solution for generating stereo image pairs.
3 Method
@@
Diverging from conventional inpainting methods, our approach is distinctively simple and training-free. With seamless integration into the original Stable Diffusion framework, we provide end-to-end generation of stereo image pairs, eliminating the need for post-processing. Our method leverages a disparity map in the early denoising stage to apply a Stereo Pixel Shift (Sec. 3.1) to the latent vector of the left image. This process generates the latent vector for the right image through disparity. To address the inconsistency issues between the left and right images during the denoising process, we employ a Symmetric Pixel Shift Masking Denoise (Sec. 3.2) technique and a Self-Attention module (Sec. 3.3) to align the right image with the left one. Since our method exclusively manipulates the latent variable, it can be applied across various image generation tasks in different Stable Diffusion models. This versatility stems from the technique’s focus on latent space operations, making it adaptable to a wide range of scenarios within the Stable Diffusion framework (Sec. 3.4). Our method only requires a disparity map, which can be obtained by various depth estimation models like DPT [26], MiDas [27] etc. and does not require camera calibration.
@@ 3.1 Stereo Pixels Shift
For the task of generating stereo images, fine-tuning models on large stereo datasets like KITTI [17] seems intuitive. However, after fine-tuning the model using various methods such as ControlNet [4] and LoRA [11], the results of the generated images remains unsatisfactory. A major flaw of this approach is that even if we could generate a high-quality stereo image pairs, the types of images generated will be limited to driving scenes similar to KITTI, losing the most important feature of Stable Diffusion: its diversity. Inspired by the Denoising Diffusion Implicit Models (DDIM) sampling technique for Stable Diffusion [32], we present a new method, Stereo Pixels Shift, without the aforementioned drawbacks. Utilizing DDIM for sampling from generalized generative processes, a latent vector sample $\bm{x}_{t-1}$ is generated from a sample $\bm{x}_{t}$ via a noise predictor $\epsilon_{\theta}$:
$\bm{x}_{t-1}=$ $\sqrt{\alpha_{t-1}}\,\underbrace{\left(\frac{\bm{x}_{t}-\sqrt{1-\alpha_{t}}\,\epsilon_{\theta}^{(t)}\!\!\left(\bm{x}_{t}\right)}{\sqrt{\alpha_{t}}}\right)}_{\text{predicted }\bm{x}_{0}}$ (1)
@@
where $\epsilon_{t}$ is noise following a standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\bm{x}_{t}$, and $\alpha_{t}$ controls the noise scale at step $t$ with $\alpha_{0}:=1$. If we set $\sigma_{t}=0$ for all $t$ and the same model $\epsilon_{\theta}$ is used, the generative results are consistent and identical, making the forward process deterministic, given $\bm{x}_{t-1}$ and $\bm{x}_{0}$. Thus, the result of $\bm{x}_{t-1}$ depends solely on $\bm{x}_{t}$. During the denoise process at a certain step $t^{\prime}$, if we modify $\bm{x}_{t^{\prime}}$ to $\bm{x}^{\prime}_{t^{\prime}}$, subsequently, $\bm{x}^{\prime}_{t^{\prime}-1}$ is denoised based on $\bm{x}^{\prime}_{t^{\prime}}$, eventually generating $\bm{x}^{\prime}_{0}$ which is different from the original $\bm{x}_{0}$. This pivotal insight enables the practical application of Stable Diffusion for stereo image generation. To align with this approach, we scale down the disparity map to match the dimensions of the latent space. Subsequently, we manipulate the latent vector on a pixel-by-pixel basis, guided by the disparity map. Given the relatively small size of the latent vector, this process does not entail a substantial computational overhead.
Assuming that the two images have parallel optical axes, we derive a disparity map from a depth map using
@@
The Stereo Pixel Shift operation $\mathcal{S}$ is expressed by
@@
The reason why we can apply Stereo Pixel Shift to the latent variable is that, after a certain step, there is a spatial position correspondence between the latent variable and the generated image. According to diffusion process theory [33, 24, 32, 44, 20, 1], sampling can be represented as
@@ 3.2 Symmetric Pixel Shift Masking Denoise
According to Eq. 1, a pixel shift applied to $x_{t^{\prime}}$ to obtain $x^{\prime}_{t^{\prime}}$ results in a slight difference between $x^{\prime}_{t^{\prime}-1}$ and $x_{t^{\prime}-1}$. This difference accumulates during the subsequent denoising process, leading to variations in the final generated image. As a result, when the denoising algorithm is applied, it may interpret the shifted areas differently, potentially causing variations in how the subject matter appears after processing. This challenge is crucial in stereo image generation, as maintaining symmetry and coherence between the two sides is essential for creating a convincing and realistic stereo effect.
To circumvent the issue, inspired by the concept of inpainting, we propose a Symmetric Pixel Shift Masking Denoise method. We create a mask for the area where the stereo pixel shift is applied. At regular intervals, defined by specific steps $t^{\prime}$, the values from the masked region of the left latent space are copied to the corresponding area of the mask in the right latent space. Consequently, the denoising process for the right image can be reformulated from Eq. 1 as
$\bm{x}^{\prime}_{t^{\prime}-1}=$ $\sqrt{\alpha_{t^{\prime}-1}}\,{\left(\frac{\bm{X}^{\prime}_{t^{\prime}}-\sqrt{1-\alpha_{t^{\prime}}}\,\epsilon_{\theta}^{(t^{\prime})}\!\left(\bm{x}^{\prime}_{t^{\prime}}\right)}{\sqrt{\alpha_{t^{\prime}}}}\right)}$ (6)
@@
where $\mathcal{S}$ represents the operation of Stereo Pixel Shift in Eq. 3, $D$ is the corresponding disparity map of the image, and $\mathbf{M}$ is a Boolean matrix of the same shape as $\bm{x}$ that signifies the mask, with values set to True for the pixels that have been shifted. The variable $\bm{x}_{t^{\prime}-1}$ denotes the latent vector of the left image at time step $t^{\prime}-1$.
@@
where $\epsilon_{t^{\prime}}$ denotes random noise, $\mathbf{M}$ is the same mask as the one in Eq. 7. However, the effectiveness varies with different images. Sometimes, it may even lead to a decrease in the quality of the generated images. A detailed effects analysis of the Deblur technique is presented in the ablation studies described in Sec. 4.3.